# Heat Flow in Semi-Flat Elliptic Fibrations

### v1

@@ 4.5 Interior bilinear heat-kernel pairing limit
$\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,(I_{t}\Phi)(x)\,(I_{t}\Psi)(y)\,d\mu_{t}(x)\,d\mu_{t}(y)\ \longrightarrow\ \int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\Phi(b)\,\Psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$ (41)
@@
4.6 Strong convergence versus bilinear pairing limits (a rigorous equivalence)
@@
$E_{t}[T_{t}(\sigma)G]\ \leq\ \frac{C}{\sigma}\,\|G\|_{L^{2}_{t}}^{2},$ (42)
@@
$\|(\Pi_{t}^{\mathrm{fc}})^{\perp}\,T_{t}(\sigma)I_{t}f\|_{L^{2}_{t}}\ \leq\ C\,\frac{s(t)}{\sqrt{\sigma}}\;\|f\|_{L^{2}_{B}},$ (43)
@@
$\|S_{t}(\tau)S_{t}(\sigma)-S_{t}(\tau+\sigma)\|_{L^{2}_{B}\to L^{2}_{B}}\ \leq\ C\,\frac{s(t)}{\sqrt{\sigma}}.$ (44)
@@
    $\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,(I_{t}\Phi)(x)\,(I_{t}\Psi)(y)\,d\mu_{t}(x)\,d\mu_{t}(y)\ \longrightarrow\ \int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\Phi(b)\,\Psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$ (45)
@@ 4.7 Section summary
5 Conic parametrix on the base and the renormalized bilinear functional
@@
5.1 Wedge charts and $C^{1}$ conic asymptotics of $g_{B}$ near $D$
@@
$g_{B}=g_{j}^{\mathrm{cone}}+q_{j},\qquad g_{j}^{\mathrm{cone}}:=dr^{2}+\alpha_{j}^{2}r^{2}\,d\theta^{2},$ (46)
@@
$|q_{j}|_{g^{\mathrm{cone}}_{j}}\leq Cr^{\beta},\qquad|\nabla^{\mathrm{cone}}q_{j}|_{g^{\mathrm{cone}}_{j}}\leq Cr^{\beta-1},$ (47)
@@
$d\mu_{g_{B}}=(1+O(r^{\beta}))\,d\mu_{g^{\mathrm{cone}}_{j}},\qquad d\mu_{g^{\mathrm{cone}}_{j}}=\alpha_{j}r\,dr\,d\theta,$ (48)
@@
5.2 The flat-cone heat kernel and basic bounds
@@
$0\leq K^{\mathrm{cone}}_{j}((r,\theta),(r^{\prime},\theta^{\prime});\tau)\leq\frac{C}{\tau}\exp\!\Big(-\frac{d_{\mathrm{cone}}((r,\theta),(r^{\prime},\theta^{\prime}))^{2}}{c\,\tau}\Big),$ (49)
@@
$|\partial_{r}K^{\mathrm{cone}}_{j}|+|\partial_{r^{\prime}}K^{\mathrm{cone}}_{j}|\leq\frac{C}{\tau^{3/2}}\exp\!\Big(-\frac{d_{\mathrm{cone}}((r,\theta),(r^{\prime},\theta^{\prime}))^{2}}{c\,\tau}\Big).$ (50)
@@
5.3 Conic parametrix for the base heat kernel with weighted remainders
@@
$K_{B}(b,b^{\prime};\tau)=\chi_{j}(b)\chi_{j}(b^{\prime})\,K^{\mathrm{cone}}_{j}(b,b^{\prime};\tau)+R_{j}(b,b^{\prime};\tau),$ (51)
@@
$|R_{j}(b,b^{\prime};\tau)|\leq\frac{C}{\tau}\,(r(b)+r(b^{\prime}))^{\beta}\,\exp\!\Big(-\frac{d_{g_{B}}(b,b^{\prime})^{2}}{C\,\tau}\Big),\qquad 0<\tau\leq\tau_{0},$ (52)
@@
5.4 Definition of the renormalized bilinear functional
@@
$K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau):=\lim_{\rho\downarrow 0}\Bigg[$ $\int_{B_{\mathrm{reg}}\times B_{\mathrm{reg}}}K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\,\Psi^{(\rho)}(b^{\prime})\,d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime})$ (54)
@@
5.5 Identification of the total-space bilinear limit with $K_{B}^{\mathrm{ren}}$
@@
$\mathcal{K}_{t}(\Phi,\Psi;\tau):=\langle I_{t}\Phi,e^{-\tau H_{t}}I_{t}\Psi\rangle_{L^{2}_{t}},\qquad\tau>0,$ (55)
@@
$\langle I_{t}\phi,e^{-\tau H_{t}}I_{t}\psi\rangle_{L^{2}(\Pi_{t}^{-1}(B(\rho)))}\longrightarrow\int_{B(\rho)\times B(\rho)}K_{B}(b,b^{\prime};\tau)\,\phi(b)\,\psi(b^{\prime})\,d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime})$ (56)
@@
$\lim_{t\downarrow 0}\mathcal{K}_{t}(\Phi,\Psi;\tau)=K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau).$ (57)
@@ 5.6 Section summary
6 Quantitative estimates and error bookkeeping
@@
6.1 Two-parameter truncation and the $\rho$–renormalized target
@@
$\mathcal{K}_{t}(\Phi,\Psi;\tau):=\langle I_{t}\Phi,\,e^{-\tau H_{t}}I_{t}\Psi\rangle_{L^{2}(X_{t})},$ (58)
@@
$K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau):=$ $\int_{B_{\mathrm{reg}}\times B_{\mathrm{reg}}}K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\,\Psi^{(\rho)}(b^{\prime})\,d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime})$ (59)
@@
6.2 Quantitative interior estimate at fixed $\rho$
@@
$\Big|\mathcal{K}_{t}(\Phi^{(\rho)},\Psi^{(\rho)};\tau)-\!\!\int_{B_{\mathrm{reg}}\times B_{\mathrm{reg}}}\!\!K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\,\Psi^{(\rho)}(b^{\prime})\,d\mu_{g_{B}}^{\otimes 2}\Big|\leq C\,\epsilon_{t}\,\|\Phi\|_{L^{2}(B)}\|\Psi\|_{L^{2}(B)}.$ (60)
@@
6.3 Mixed-term control
@@
$\|\Phi^{<\rho}\|_{L^{2}(B)}\leq C\,\rho\,\|\Phi\|_{L^{\infty}(B)},\qquad\|\Psi^{<\rho}\|_{L^{2}(B)}\leq C\,\rho\,\|\Psi\|_{L^{\infty}(B)}.$ (63)
@@
6.4 Edge parametrix remainder: the $O(\rho^{4+\beta})$ term
@@
$\big|K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)\big|\leq C\,\rho^{4+\beta}\,\|\Phi\|_{L^{\infty}(B)}\,\|\Psi\|_{L^{\infty}(B)},\qquad\tau\in[\tau_{1},\tau_{2}].$ (64)
@@
6.5 Final bookkeeping and the citable iterated-limit theorem
@@
$\Big|\mathcal{K}_{t}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau)\Big|\leq\;$ $C(\rho,\tau_{1},\tau_{2})\,\epsilon_{t}\,\|\Phi\|_{L^{2}(B)}\|\Psi\|_{L^{2}(B)}$ (65)
@@
$\lim_{\rho\downarrow 0}\ \limsup_{t\downarrow 0}\ \Big|\mathcal{K}_{t}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)\Big|=0,\qquad\tau\in[\tau_{1},\tau_{2}].$ (66)
@@
7 Synthesis: renormalized limit for compressed heat flow
@@
$\mathcal{K}_{t}(\Phi,\Psi;\tau):=\langle I_{t}\Phi,\,e^{-\tau H_{t}}I_{t}\Psi\rangle_{L^{2}(X_{t})}\longrightarrow K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)\qquad(t\downarrow 0).$ (67)
@@
$\big|\mathcal{K}_{t}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau)\big|\leq\mathrm{Err}_{\mathrm{int}}(t;\Phi,\Psi)+\mathrm{Err}_{\mathrm{mix}}(\rho,\tau;\Phi,\Psi)+\mathrm{Err}_{\mathrm{edge}}(\rho,\tau;\Phi,\Psi),$ (68)
@@
8 Geometric input in semi-flat elliptic fibrations

### v2

@@ 4.5. Interior bilinear heat-kernel pairing limit
(42) $\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,(I_{t}\Phi)(x)\,(I_{t}\Psi)(y)\,d\mu_{t}(x)\,d\mu_{t}(y)\ \longrightarrow\ \int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\Phi(b)\,\Psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$
@@
4.6. Strong convergence versus bilinear pairing limits (a rigorous equivalence)
@@
(43) $E_{t}[T_{t}(\sigma)G]\ \leq\ \frac{C}{\sigma}\,\|G\|_{L^{2}_{t}}^{2},$
@@
(44) $\|(\Pi_{t}^{\mathrm{fc}})^{\perp}\,T_{t}(\sigma)I_{t}f\|_{L^{2}_{t}}\ \leq\ C\,\frac{s(t)}{\sqrt{\sigma}}\;\|f\|_{L^{2}_{B}},$
@@
(45) $\|S_{t}(\tau)S_{t}(\sigma)-S_{t}(\tau+\sigma)\|_{L^{2}_{B}\to L^{2}_{B}}\ \leq\ C\,\frac{s(t)}{\sqrt{\sigma}}.$
@@
    (46) $\int_{X_{t,0}\times X_{t,0}}K_{t}(x,y;\tau)\,(I_{t}\Phi)(x)\,(I_{t}\Psi)(y)\,d\mu_{t}(x)\,d\mu_{t}(y)\ \longrightarrow\ \int_{B_{0}\times B_{0}}K_{B}(b,b^{\prime};\tau)\,\Phi(b)\,\Psi(b^{\prime})\,d\mu_{B}(b)\,d\mu_{B}(b^{\prime}).$
@@ 4.7. Section summary
5. Conic parametrix on the base and the renormalized bilinear functional
@@
5.1. Wedge charts and $C^{1}$ conic asymptotics of $g_{B}$ near $D$
@@
(47) $g_{B}=g_{j}^{\mathrm{cone}}+q_{j},\qquad g_{j}^{\mathrm{cone}}:=dr^{2}+\alpha_{j}^{2}r^{2}\,d\theta^{2},$
@@
(48) $|q_{j}|_{g^{\mathrm{cone}}_{j}}\leq Cr^{\beta},\qquad|\nabla^{\mathrm{cone}}q_{j}|_{g^{\mathrm{cone}}_{j}}\leq Cr^{\beta-1},$
@@
(49) $d\mu_{g_{B}}=(1+O(r^{\beta}))\,d\mu_{g^{\mathrm{cone}}_{j}},\qquad d\mu_{g^{\mathrm{cone}}_{j}}=\alpha_{j}r\,dr\,d\theta,$
@@
See Figure 2 for the wedge chart and the inner/outer decomposition used in the renormalization.
[…]
Figure 2. Local wedge chart near a discriminant point $p_{j}\in D$ and the corresponding flat-cone model. Panel (a) shows polar coordinates $(r,\theta)$ on $U_{j}\setminus\{p_{j}\}\subset B_{\mathrm{reg}}$, together with the inner region $r<\rho$ and outer region $r>\rho$ used in the renormalization. Panel (b) depicts the model cone $\mathsf{C}_{\alpha_{j}}$ with cone angle $2\pi\alpha_{j}$, for which $g_{B}$ is $C^{1}$-close in Assumption 5.1.
5.2. The flat-cone heat kernel and basic bounds
@@
(50) $0\leq K^{\mathrm{cone}}_{j}((r,\theta),(r^{\prime},\theta^{\prime});\tau)\leq\frac{C}{\tau}\exp\!\Big(-\frac{d_{\mathrm{cone}}((r,\theta),(r^{\prime},\theta^{\prime}))^{2}}{c\,\tau}\Big),$
@@
(51) $|\partial_{r}K^{\mathrm{cone}}_{j}|+|\partial_{r^{\prime}}K^{\mathrm{cone}}_{j}|\leq\frac{C}{\tau^{3/2}}\exp\!\Big(-\frac{d_{\mathrm{cone}}((r,\theta),(r^{\prime},\theta^{\prime}))^{2}}{c\,\tau}\Big).$
@@
5.3. Conic parametrix for the base heat kernel with weighted remainders
@@
(52) $K_{B}(b,b^{\prime};\tau)=\chi_{j}(b)\chi_{j}(b^{\prime})\,K^{\mathrm{cone}}_{j}(b,b^{\prime};\tau)+R_{j}(b,b^{\prime};\tau),$
@@
(53) $|R_{j}(b,b^{\prime};\tau)|\leq\frac{C}{\tau}\,(r(b)+r(b^{\prime}))^{\beta}\,\exp\!\Big(-\frac{d_{g_{B}}(b,b^{\prime})^{2}}{C\,\tau}\Big),\qquad 0<\tau\leq\tau_{0},$
@@
5.4. Definition of the renormalized bilinear functional
@@
(55) $K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau):=\lim_{\rho\downarrow 0}\Bigg[$ $\int_{B_{\mathrm{reg}}\times B_{\mathrm{reg}}}K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\,\Psi^{(\rho)}(b^{\prime})\,d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime})$
@@
[…]
$\Phi=\Phi^{(\rho)}+\Phi^{<\rho},\qquad\Psi=\Psi^{(\rho)}+\Psi^{<\rho}$
Figure 3. $\rho$–cutoff decomposition and the gluing mechanism in the definition of the renormalized functional $K_{B}^{\mathrm{ren}}$. The test functions split into an outer part supported in $B_{\mathrm{reg}}$ and an inner part supported near the conic points. The outer contribution uses the base heat kernel $K_{B}$, while the inner contribution is replaced by the model cone kernels $K_{j}^{\mathrm{cone}}$. Taking $\rho\downarrow 0$ yields a canonical cutoff-independent limit (Theorem 5.7).
@@
5.5. Identification of the total-space bilinear limit with $K_{B}^{\mathrm{ren}}$
@@
(56) $\mathcal{K}_{t}(\Phi,\Psi;\tau):=\langle I_{t}\Phi,e^{-\tau H_{t}}I_{t}\Psi\rangle_{L^{2}_{t}},\qquad\tau>0,$
@@
(57) $\langle I_{t}\phi,e^{-\tau H_{t}}I_{t}\psi\rangle_{L^{2}(\Pi_{t}^{-1}(B(\rho)))}\longrightarrow\int_{B(\rho)\times B(\rho)}K_{B}(b,b^{\prime};\tau)\,\phi(b)\,\psi(b^{\prime})\,d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime})$
@@
(58) $\lim_{t\downarrow 0}\mathcal{K}_{t}(\Phi,\Psi;\tau)=K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau).$
@@ 5.6. Section summary
6. Quantitative estimates and error bookkeeping
@@
6.1. Two-parameter truncation and the $\rho$–renormalized target
@@
(59) $\mathcal{K}_{t}(\Phi,\Psi;\tau):=\langle I_{t}\Phi,\,e^{-\tau H_{t}}I_{t}\Psi\rangle_{L^{2}(X_{t})},$
@@
(60) $K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau):=$ $\int_{B_{\mathrm{reg}}\times B_{\mathrm{reg}}}K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\,\Psi^{(\rho)}(b^{\prime})\,d\mu_{g_{B}}(b)\,d\mu_{g_{B}}(b^{\prime})$
@@
6.2. Quantitative interior estimate at fixed $\rho$
@@
(61) $\Big|\mathcal{K}_{t}(\Phi^{(\rho)},\Psi^{(\rho)};\tau)-\!\!\int_{B_{\mathrm{reg}}\times B_{\mathrm{reg}}}\!\!K_{B}(b,b^{\prime};\tau)\,\Phi^{(\rho)}(b)\,\Psi^{(\rho)}(b^{\prime})\,d\mu_{g_{B}}^{\otimes 2}\Big|\leq C\,\epsilon_{t}\,\|\Phi\|_{L^{2}(B)}\|\Psi\|_{L^{2}(B)}.$
@@
6.3. Mixed-term control
@@
(64) $\|\Phi^{<\rho}\|_{L^{2}(B)}\leq C\,\rho\,\|\Phi\|_{L^{\infty}(B)},\qquad\|\Psi^{<\rho}\|_{L^{2}(B)}\leq C\,\rho\,\|\Psi\|_{L^{\infty}(B)}.$
@@
6.4. Edge parametrix remainder: the $O(\rho^{4+\beta})$ term
@@
(65) $\big|K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)\big|\leq C\,\rho^{4+\beta}\,\|\Phi\|_{L^{\infty}(B)}\,\|\Psi\|_{L^{\infty}(B)},\qquad\tau\in[\tau_{1},\tau_{2}].$
@@
6.5. Final bookkeeping and the citable iterated-limit theorem
@@
(66) $\Big|\mathcal{K}_{t}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau)\Big|\leq\;$ $C(\rho,\tau_{1},\tau_{2})\,\epsilon_{t}\,\|\Phi\|_{L^{2}(B)}\|\Psi\|_{L^{2}(B)}$
@@
(67) $\lim_{\rho\downarrow 0}\ \limsup_{t\downarrow 0}\ \Big|\mathcal{K}_{t}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B}(\Phi,\Psi;\tau)\Big|=0,\qquad\tau\in[\tau_{1},\tau_{2}].$
@@
7. Synthesis: renormalized limit for compressed heat flow
@@
(68) $\mathcal{K}_{t}(\Phi,\Psi;\tau):=\langle I_{t}\Phi,\,e^{-\tau H_{t}}I_{t}\Psi\rangle_{L^{2}(X_{t})}\longrightarrow K_{B}^{\mathrm{ren}}(\Phi,\Psi;\tau)\qquad(t\downarrow 0).$
@@
(69) $\big|\mathcal{K}_{t}(\Phi,\Psi;\tau)-K^{\mathrm{ren}}_{B,\rho}(\Phi,\Psi;\tau)\big|\leq\mathrm{Err}_{\mathrm{int}}(t;\Phi,\Psi)+\mathrm{Err}_{\mathrm{mix}}(\rho,\tau;\Phi,\Psi)+\mathrm{Err}_{\mathrm{edge}}(\rho,\tau;\Phi,\Psi),$
@@
8. Geometric input in semi-flat elliptic fibrations